# Activated solids: Spontaneous deformations, non-affine fluctuations, softening, and failure

### v2

@@ I Introduction
The study of active solids [1, 2, 3, 4, 5, 6] is inspired by living systems such as tissues and cell assemblies [7, 8, 9], bacterial biofilms [10, 11], and artificial material like smart matter [12, 13]. These systems are inherently out of equilibrium at the microscopic scale: their constituents locally consume and dissipate energy, generating stresses or propulsion and breaking detailed balance. Unlike crystalline solids, which break continuous translational and rotational symmetries in equilibrium, amorphous solids are stabilized by compression and by local force and torque balance [14, 15, 16, 17]. In active solids maintaining this balance is essential for mechanical stability; its breakdown naturally lead to deformation and flow.
@@
We investigate spontaneous deformations in active solids by quantifying local non-affine fluctuations relative to the global strain. Our main achievements are threefold. First, using scaling arguments and numerical simulations, we predict how the steady-state mean non-affinity depends on activity parameters – active speed and persistence – revealing behavior beyond simple active-temperature estimates. Second, at higher activity, pronounced non-affine zones coexist with regions of small non-affinity, preceding the divergence of non-affinity, defect proliferation, and eventual melting. Correlation functions show that the non-affine correlation length increases with active speed and grows to saturation with active persistence, while the relaxation of system-averaged non-affinity is set by the active persistence time. Third, we demonstrate that spatially patterned activation provides a direct route to locally induce non-affinity and mechanical softening. Together, these results establish a quantitative framework for understanding and controlling the interplay between activity, non-affine deformations, and mechanical stability in dense active materials.
@@ II Model
dθ_i(t) = 2D_r dB^r_i(t), where $D_{t}$ and $D_{r}$ are the translational and rotational diffusivities, and $d\mathbf{B}_{i}(t)$ and $dB_{i}^{r}(t)$ are independent Gaussian white noise terms with mean zero and variance $dt$. The Euler-Maruyama method is used to directly integrate the dimensionless forms of the above equations (see Appendix-A) in numerical simulations. The units of energy, length, and time are set by $\epsilon$, $\sigma$, $\tau_{u}=\sigma^{2}/\mu\epsilon$. The system property is controlled by the following dimensionless parameters: active speed $\Lambda=v_{0}\sigma/\mu\epsilon$, translational diffusivity $\tilde{D}_{t}=D_{t}/\mu\epsilon$, and orientational diffusivity $\tilde{D}_{r}=D_{r}\sigma^{2}/\mu\epsilon$. Unless stated otherwise, we set $\tilde{D}_{t}=1$ and $\tilde{D}_{t}=0$, to simulate thermal or athermal systems, respectively, and fix the density at $\rho\sigma^{2}=1.1$. We also examine the effects of activity by varying $\Lambda$} and $\tilde{D}_{r}$, as well as the effect of changing density $\rho\sigma^{2}$.
@@ III Non-affine parameter
Figure 1: A schematic illustration of the activated solid. (a) An ideal triangular lattice that serves as the reference for the calculation of non-affinity $\chi$. (b) A representative configuration at $\tilde{D}_{t}=1.0$ and $\Lambda=8.0$, is shown color coded by local non-affine parameter $\chi$; see the color bar. (c) A single coarse-graining region $\Omega$ is defined as the volume enclosed by the first nearest neighbors, determined by mean inter-particle separation $a$ at a given denisty $\rho$. This region includes one central particle with reference position $\mathbf{R}_{0}$ (cyan) and actual location $\mathbf{r}_{0}$ (dark green). The reference (cyan, $\{\mathbf{R}_{i}\}$) and actual locations (dark green, $\{\mathbf{r}_{i}\}$) of its six nearest neighbors are shown as well.
@@ III.1 Scaling analysis
In an equilibrium solid with harmonic relaxation rate $\mu k$ and single particle diffusivity $D_{t}$, the characteristic length scale associated with particle rearrangements, $\ell_{\rm eq}$, is set by a balance of diffusion and relaxation, yielding $\ell_{\rm eq}^{2}\sim D_{t}/\mu k$. For an equilibrium harmonic solid, the $n$-th moment of the mean non-affine parameter, $X=N^{-1}\sum_{i}^{N}\chi_{i}$, was shown to scale as $\langle X^{n}\rangle_{eq}\sim\ell^{2n}$ [31]. In sterically stabilized solids, the effective elastic constant $k$ is determined by $U^{\prime\prime}(a)$, and therefore increases with density. For a two-dimensional isotropic solid, the shear and Young’s moduli scale linearly with $k$ and, more generally, increase nonlinearly with density [35]. This elastic stiffening implies a corresponding decrease in the mean non-affinity, $\langle X\rangle$. Approaching melting, however, the shear modulus $\cal G$} vanishes, and the non-affine parameter is expected to diverge. These considerations motivate a generalized expression for equilibrium non-affine rearrangements, $\langle X\rangle_{eq}\sim\frac{D_{t}}{{\cal G}}$, with $\cal G$} increasing with density $\rho$ and vanishing at the melting transition. Ignoring non-linearities for simplicity, the density dependence of shear modulus can be approximated as:
${\cal G}(\rho)$ $=$ $a(\rho-\rho_{m})\quad\text{for}\quad\rho>\rho_{m},$ (4)
@@
[…]
In this paper, we first consider the impact of activity on a dense solid. As activity increases, the persistence length $\ell_{p}=v_{0}/D_{r}$ begins to play an important role. One might then expect that replacing the thermal diffusivity $D_{t}$ by the effective active diffusivity $v_{0}^{2}/2D_{r}$ [36, 37] in the expression $D_{t}/{\cal G}$ would be sufficient to account for the excess non-affinity induced by activity. However, this simple substitution fails in a solid, where each test particle is trapped within a cage formed by its neighbors. Approximating this confinement by a harmonic potential with stiffness $k$, we can invoke an earlier estimate for the steady-state mean-squared displacement arising from activity $\ell_{\rm ac}^{2}=v_{0}^{2}/[\mu k(D_{r}+\mu k)]$, as obtained in Ref. [38]. This leads us to hypothesize that activity induces an excess non-affinity of the form $\langle X^{(a)}\rangle\approx v_{0}^{2}/[\alpha{\cal G}(D_{r}+\alpha{\cal G})]$, where the spring constant is replaced by the more appropriate solid descriptor, the shear modulus $\cal G$}; here $\alpha$ is a proportionality constant. Thus, the total non-affine parameter can be written in the following dimensionless form as
@@
In summary, for densities $\rho>\rho_{m}$ and active speed $\Lambda<\Lambda_{c}$ allowing the system to remain in the solid phase, the following scaling forms for $\langle X^{(a)}\rangle$ are expected:
$\langle X^{(a)}\rangle\sim(\rho-\rho_{m})^{-1}\quad\text{and}\quad\langle X^{(a)}\rangle\sim\frac{\Lambda^{2}}{\tilde{\cal G}[\tilde{D}_{r}+\alpha\tilde{\cal G}]}$. (7)
@@ III.2 Scaling verification via numerical simulations
In Fig. 2(c), we present simulations of $N=4096$ ABPs for $\tilde{D}_{r}=3.0$, $\Lambda\in[0.5,8.0]$, densities $\rho\in[1.0,1.4]$, and $\tilde{D}_{t}=1.0$, from which we extract the excess steady-state non-affinity due to activity $\langle X^{(a)}\rangle$. These results are compared with athermal simulations performed at $\tilde{D}_{t}=0$. A reasonable data collapse is obtained when results at different densities are plotted as $\langle X^{(a)}\rangle/a_{\rho}$ versus $\Lambda$. The excess non-affinity follows a scaling form $\langle X^{(a)}\rangle/a_{\rho}\sim\Lambda^{\nu}$ with exponent $\nu\approx 2.0$, as shown in Fig. 2(c). The inset of Fig. 2(c) displays the density dependence of the prefactor, $a_{\rho}\sim(\rho-\rho_{0})^{-1}$, with $\rho_{0}\sigma^{2}=0.971\pm 0.004$. This fitted threshold exceeds the equilibrium melting density $\rho_{m}^{\mathrm{eq}}\sigma^{2}=0.92$ [34], indicating activity-induced softening of the solid.
@@ III.3 Coexistence
In Fig.3, we show the probability distributions $P(\chi)$ of local non-affinity $\chi_{i}=\chi$ for an athermal system ($\tilde{D}_{t}=0$) on log-log axes for varying active speed (panel (a) ) and persistence time (panel (b) ). In equilibrium, the distribution is described by a chi-squared form [41]. As the activity parameters increase, the slightly skewed distributions broaden, get more skewed, and develop a heavy tail, followed by the appearance of a secondary maximum indicating coexisting small and large non-affinities. Such coexistence of pronounced non-affine zones with an otherwise normal solid precedes the divergence of non-affinity, which emerges as a precursor to defect formation and eventual melting.
@@ III.4 Correlations
at steady state, where $r=|{\mathbf{r}}|$, $\chi_{i}$ denotes the local non-affinity for particle $i$, and $r_{ij}=|{\mathbf{r}}_{i}-{\mathbf{r}}_{j}|$ is the distance between particles $i$ and $j$. The normalized two-point correlation function $C\chi(|\mathbf{r}|)/C\chi(0)$ at fixed activity $\Lambda=5$ is shown in Fig. 4(a). These correlations clearly extend over longer distances as persistence increases (i.e., as $\tilde{D}_{r}^{-1}$ grows). We extract a characteristic correlation length $\xi$ using the integral definition $\xi=\int dr\,C_{\chi}(r)/C_{\chi}(0)$, which captures the correct trends, although the precise numerical value can vary depending on the specific choice of correlation length measure. Mapping $\xi$ across a range of $\Lambda$ and $\tilde{D}_{r}$ (Fig. 4(b)) reveals three distinct regimes. For high persistence, $\tilde{D}_{r}\lesssim 0.2$, $\xi$ saturates to a $\Lambda$-dependent plateau, reflecting extended correlations in persistent-active systems. In the intermediate regime, $0.2\lesssim\tilde{D}_{r}\lesssim 110$, $\xi$ decreases monotonically with $\tilde{D}_{r}$, indicating progressively shorter-ranged correlations as persistence weakens. Finally, in the low-persistence regime, $\tilde{D}_{r}\gtrsim 110$, the short correlation lengths become largely independent of activity. Increasing $\Lambda$ systematically enhances $\xi$ in the persistent-active and intermediate regimes, showing that stronger self-propulsion promotes correlations of non-affine rearrangements over longer distances, whereas in the low-persistence limit the correlation length collapses to a small value, independent of $\Lambda$.
@@ IV Analyzing structural indicators
Thus far, we have characterized the softening of a high-density solid through the non-affine parameter $X$ and its scaling with activity characterized by persistence and active speed, and density. In this section, we extend our analysis to include elastic and structural indicators, such as the shear modulus, solid order, and hexatic order, to further explore the active softening of solid. We perform this analysis using $\rho\sigma^{2}=1.1$, $\tilde{D}_{r}=3.0$ and $\tilde{D}_{t}=1.0$.
@@ IV.1 Shear modulus
Within linear response, the shear modulus is given by $\mathcal{G}=\sigma^{tot}_{xy}/{\cal E}_{xy}$. We compute this from numerical simulations by applying a shear strain ${\cal E}_{xy}$ and measuring the change in total shear stress $\sigma_{xy}^{\mathrm{tot}}$ accounting for trajectories displaying linear response; see Appendix-E. The stress tensor has contribution from interaction calculated using the usual virial term and swim stress due to activity. The standard virial expression due to interaction gives,
@@
The total shear stress is then calculated as $\sigma^{tot}_{xy}=\sigma^{vir}_{xy}+\sigma^{sw}_{xy}$. Appendix-E outlines details of the shear modulus calculation.
Figure 6(a) shows the change in mean shear modulus, $\cal G$, with standard errors, as a function of activity $\Lambda$. We observe a steady decrease in it from the equilibrium value indicating softening of the solid with activity. As activity increases, fraction of trajectories following linear response decreases to vanish near $\Lambda=10$; see Appendix-E. The decrease in $G$ with activity can be described by ${\cal G}={\cal G}_{0}-{\cal G}_{1}\Lambda^{2}$ with ${\cal G}_{0}=109.84$ and ${\cal G}_{1}=0.094$. To further explore this softening and the fluidization transition, we compute structural indicators, including solid order $\psi_{\mathbf{G}}$, hexatic order $\psi_{6}$, and defect fraction $\delta_{d}$.

### v3

@@ I Introduction
The study of active solids [1, 2, 3, 4, 5, 6] is inspired by living systems such as tissues and cell assemblies [7, 8, 9], bacterial biofilms [10, 11], and artificial material such as smart matter [12, 13]. These systems are inherently out of equilibrium at the microscopic scale: their constituents locally consume and dissipate energy, generating stresses or propulsion and breaking detailed balance. Unlike crystalline solids, which break continuous translational and rotational symmetries in equilibrium, amorphous solids are stabilized by compression and by local force and torque balance [14, 15, 16, 17]. In active solids, maintaining this balance is essential for mechanical stability; its breakdown naturally leads to deformation and flow.
@@
We investigate spontaneous deformations in active solids by quantifying local non-affine fluctuations relative to the global strain. Our main achievements are threefold. First, using scaling arguments and numerical simulations, we predict how the steady-state mean non-affinity depends on activity parameters – active speed and persistence – revealing behavior beyond simple active-temperature estimates. Second, at higher activity, pronounced non-affine zones coexist with regions of small non-affinity, preceding the divergence of non-affinity, defect proliferation, and eventual melting. Correlation functions show that the non-affine correlation length increases with active speed and grows to finally saturate with active persistence, while the relaxation of system-averaged non-affinity is set by the active persistence time. Third, we demonstrate that spatially patterned activation provides a direct route to locally induce non-affinity and mechanical softening. Together, these results establish a quantitative framework for understanding and controlling the interplay between activity, non-affine deformations, and mechanical stability in dense active materials.
@@ II Model
dθ_i(t) = 2D_r dB^r_i(t), where $D_{t}$ and $D_{r}$ are the translational and rotational diffusivities, and $d\mathbf{B}_{i}(t)$ and $dB_{i}^{r}(t)$ are independent Gaussian white noise terms with mean zero and variance $dt$. The Euler-Maruyama method is used to directly integrate the dimensionless forms of the above equations (see Appendix-A) in numerical simulations. The units of energy, length, and time are set by $\epsilon$, $\sigma$, $\tau_{u}=\sigma^{2}/\mu\epsilon$. The system property is controlled by the following dimensionless parameters: active speed $\Lambda=v_{0}\sigma/\mu\epsilon$, translational diffusivity $\tilde{D}_{t}=D_{t}/\mu\epsilon$, and orientational diffusivity $\tilde{D}_{r}=D_{r}\sigma^{2}/\mu\epsilon$. Unless stated otherwise, we set $\tilde{D}_{t}=1$ for thermal systems and $\tilde{D}_{t}=0$ for athermal systems, and fix the density at $\rho\sigma^{2}=1.1$. We examine the effects of activity by varying $\Lambda$} and $\tilde{D}_{r}$, and additionally investigate the effect of changing density $\rho\sigma^{2}$ in one case.
@@ III Non-affine parameter
Figure 1: A schematic illustration of the activated solid. (a) An ideal triangular lattice that serves as the reference for the calculation of non-affinity $\chi$. (b) A representative configuration at $\tilde{D}_{t}=1.0$ and $\Lambda=8.0$, is shown color coded by local non-affine parameter $\chi$; see the color bar. (c) A single coarse-graining region $\Omega$ is defined as the volume enclosed by the first nearest neighbors, determined by mean inter-particle separation $a$ at a given density $\rho$. This region includes one central particle with reference position $\mathbf{R}_{0}$ (cyan) and actual location $\mathbf{r}_{0}$ (dark green). The reference (cyan, $\{\mathbf{R}_{i}\}$) and actual locations (dark green, $\{\mathbf{r}_{i}\}$) of its six nearest neighbors are shown as well.
@@ III.1 Scaling analysis
The average total non-affinity, $\langle\sum_{i}\chi_{i}\rangle=\mathrm{Tr}[PC]$ [31], with $C=\langle\Delta\Delta^{T}\rangle$ and $\Delta_{i}={\bf u}_{i}-{\bf u}_{0}$, connects non-affinity to relative displacements. The projection operators P=IQP=I-Q and $Q=R(R^{T}R)^{-1}R^{T}$ are symmetric and idempotent projection operators, projecting onto non-affine and affine subspaces, respectively, with $R$ a block matrix of identities and zeros determined by lattice geometry. Consequently, the mean global non-affinity, $\langle X\rangle$, scales with $\langle C\rangle\sim\ell^{2}$, directly linking non-affinity to $\ell^{2}$, the mean-squared displacement of particles  [^1] More precisely, $\ell^{2}$ corresponds to the mean-squared displacement of the normal modes.
In an equilibrium solid with harmonic interactions of stiffness $k$, the single-particle diffusivity $D_{t}$ and the harmonic relaxation rate $\mu k$ set the mean-squared displacement, $\ell_{\rm eq}^{2}\sim D_{t}/(\mu k)$. For an equilibrium harmonic solid, the $n$-th moment of $X$, was shown to scale as $\langle X^{n}\rangle_{eq}\sim\ell_{eq}^{2n}$ [31]. In sterically stabilized solids, $k$ is determined by $U^{\prime\prime}(a)$, and therefore increases with density. For a two-dimensional isotropic solid, the shear and Young’s moduli scale linearly with $k$ and, more generally, increase nonlinearly with density [36]. This elastic stiffening implies a corresponding decrease in the mean non-affinity, $\langle X\rangle$. Approaching melting, however, the shear modulus $\cal G$} vanishes, and the non-affine parameter is expected to diverge. These considerations motivate a generalized expression for equilibrium non-affine rearrangements, $\langle X\rangle_{eq}\sim\frac{D_{t}}{{\cal G}}$, with $\cal G$} increasing with density $\rho$ and vanishing at the melting transition. Ignoring non-linearities for simplicity, the density dependence of shear modulus can be approximated as:
${\cal G}(\rho)$ $\propto$ $(\rho-\rho_{m})\quad\text{for}\quad\rho>\rho_{m},$ (4)
@@
with $\rho_{m}$ representing the melting density. As shown below, this hypothesis predicts that non-affinity decreases with increasing density, a trend confirmed by numerical simulations.
In this paper, we consider the impact of activity on a dense solid. One might expect that replacing the thermal diffusivity $D_{t}$ with the effective active diffusivity $v_{0}^{2}/2D_{r}$ [37, 38] in $D_{t}/{\cal G}$ would account for the activity-induced excess non-affinity. However, this simple substitution fails in a solid, where each test particle is trapped within a cage formed by its neighbors. Approximating this confinement by a harmonic potential with stiffness $k$, we can invoke an earlier estimate for the steady-state mean-squared displacement arising from activity $\ell_{\rm ac}^{2}=v_{0}^{2}/[\mu k(D_{r}+\mu k)]$, as obtained in Ref. [39]. This leads us to hypothesize that activity induces an excess non-affinity of the form $\langle X^{(a)}\rangle\approx v_{0}^{2}/[\alpha{\cal G}(D_{r}+\alpha{\cal G})]$, where the spring constant is replaced by the more appropriate solid descriptor, the shear modulus $\cal G$}; here $\alpha$ is a proportionality constant. Thus, the total non-affine parameter can be written in the following dimensionless form as
@@
In summary, for densities $\rho>\rho_{m}$ and active speed $\Lambda<\Lambda_{c}$ allowing the system to remain in the solid phase, the following scaling form for $\langle X^{(a)}\rangle$ is expected:
$\langle X^{(a)}\rangle\sim\frac{\Lambda^{2}}{\tilde{\cal G}[\tilde{D}_{r}+\alpha\tilde{\cal G}]}$. (7)
@@ III.2 Scaling verification via numerical simulations
In Fig. 2(c), we present simulations of $N=4096$ ABPs for $\tilde{D}_{r}=3.0$, $\Lambda\in[0.5,8.0]$, densities $\rho\in[1.0,1.4]$, and $\tilde{D}_{t}=1.0$, from which we extract the excess steady-state non-affinity due to activity $\langle X^{(a)}\rangle$. These results are compared with athermal simulations performed at $\tilde{D}_{t}=0$. A reasonable data collapse is obtained when results at different densities are plotted as $\langle X^{(a)}\rangle/a_{\rho}$ versus $\Lambda$. The excess non-affinity follows a scaling form $\langle X^{(a)}\rangle/a_{\rho}\sim\Lambda^{\nu}$ with exponent $\nu\approx 2.0$, as shown in Fig. 2(c).
@@
[…]
The scaling of global non-affinity with activity parameters reflects the mechanisms by which activity drives particle rearrangements in solids. Increasing the self-propulsion speed, $\Lambda$, strengthens active fluctuations and thereby enhances non-affinity. Increasing the persistence time, $\tilde{D}_{r}^{-1}$, allows particles to keep their heading directions unchanged for longer durations, promoting more coherent local rearrangements. Consequently, $\langle X\rangle$ increases with persistence at low to moderate values, but saturates at large persistence as the system gets effectively jammed. This behavior arises from the competition between persistent active forcing and structural constraints of the solid, and contrasts with equilibrium solids, where particle displacements are isotropic, homogeneous, and Gaussian.
@@ III.3 Coexistence
In Fig.3, we show the steady-state probability distributions $P(\chi)$ of local non-affinity $\chi_{i}=\chi$ for an athermal system ($\tilde{D}_{t}=0$) on log-log axes for varying active speed (panel (a) ) and persistence time (panel (b) ). In equilibrium, the distribution is described by a chi-squared form [43], reflecting the Gaussian statistics of particle displacements and the quadratic dependence of the local non-affinity on these fluctuations. We observe a clear departure from this expectation in the spontaneous fluctuations of active solids, consistent with persistence rendering non-Gaussianity to displacement fluctuations [44]. As the activity parameters increase, the slightly skewed distributions broaden, get more skewed, and develop a heavy tail, followed by the appearance of a secondary maximum indicating rare but large non-affine rearrangements. This bimodality, arising from pronounced non-affine zones within an otherwise normal solid, precedes the divergence of non-affinity and serves as a precursor to defect formation and eventual melting. The statistical weight of large non-affinity originates from either transient formation and decay of highly non-affine regions or steady-state coexistence with local plastic zones; both types of events are illustrated in Appendix-E using particle configurations and time series of the global non-affinity.
@@ III.4 Correlations
at steady state, where $r=|{\mathbf{r}}|$, $\chi_{i}$ denotes the local non-affinity for particle $i$, and $r_{ij}=|{\mathbf{r}}_{i}-{\mathbf{r}}_{j}|$ is the distance between particles $i$ and $j$. The normalized two-point correlation function $C_{\chi}(|\mathbf{r}|)/C_{\chi}(0)$ at fixed activity $\Lambda=5$ is shown in Fig. 4(a). These correlations clearly extend over longer distances as persistence increases (i.e., as $\tilde{D}_{r}^{-1}$ grows). We extract a characteristic correlation length $\xi$ using the integral definition $\xi=\int dr\,C_{\chi}(r)/C_{\chi}(0)$, which captures the correct trends, although the precise numerical value can vary depending on the specific choice of correlation length measure. Mapping $\xi$ across a range of $\Lambda$ and $\tilde{D}_{r}$ (Fig. 4(b)) reveals three distinct regimes. For high persistence, $\tilde{D}_{r}\lesssim 0.2$, $\xi$ saturates to a $\Lambda$-dependent plateau, reflecting extended correlations in persistent-active systems. In the intermediate regime, $0.2\lesssim\tilde{D}_{r}\lesssim 110$, $\xi$ decreases monotonically with $\tilde{D}_{r}$, indicating progressively shorter-ranged correlations as persistence weakens. Finally, in the low-persistence regime, $\tilde{D}_{r}\gtrsim 110$, the short correlation lengths become largely independent of activity. Increasing $\Lambda$ systematically enhances $\xi$ in the persistent-active and intermediate regimes, showing that stronger self-propulsion promotes correlations of non-affine rearrangements over longer distances, whereas in the low-persistence limit the correlation length collapses to a small value, independent of $\Lambda$.
@@
Active solids display behavior that is fundamentally different from their equilibrium counterparts. In equilibrium systems, particle displacements are governed by Gaussian fluctuations, leading to chi-squared distributions of local non-affinity and correlations that decay within roughly two lattice spacings [43, 50]. Active particles, however, apply persistently directed forces along their instantaneous headings. While at low persistence they behave similarly to passive particles, increasing persistence enhances directional, non-Gaussian fluctuations. As a result, the distributions of local non-affinity deviate strongly from the chi-squared form (Fig. 3), and the associated correlation length grows with persistence (Fig. 4), indicating the emergence of longer-ranged cooperative rearrangements absent in equilibrium solids.
@@ IV Analyzing structural indicators
Thus far, we have characterized the softening of a high-density solid through the non-affine parameter $X$ and its scaling with activity characterized by persistence and active speed, and density. In this section, we extend our analysis to include elastic and structural indicators, such as the shear modulus, solid order, and hexatic order, to further explore the active softening of the solid. We perform this analysis using $\rho\sigma^{2}=1.1$, $\tilde{D}_{r}=3.0$ and $\tilde{D}_{t}=1.0$.
@@ IV.1 Shear modulus
Within linear response, the shear modulus is given by $\mathcal{G}=\sigma^{tot}_{xy}/{\cal E}_{xy}$. We compute this from numerical simulations by applying a shear strain ${\cal E}_{xy}$ and measuring the change in total shear stress $\sigma_{xy}^{\mathrm{tot}}$ accounting for trajectories displaying linear response; see Appendix-F. The stress tensor has contributions from interaction calculated using the usual virial term and swim stress due to activity. The standard virial expression due to interaction gives,
@@
The total shear stress is then calculated as $\sigma^{tot}_{xy}=\sigma^{vir}_{xy}+\sigma^{sw}_{xy}$. Appendix-F outlines details of the shear modulus calculation. The mean swim stress does not change under shear, as confirmed by numerical calculations; so the shear modulus is set solely by the virial stress.
Figure 6(a) shows the change in mean shear modulus, $\cal G$, with standard errors, as a function of activity $\Lambda$. We observe a steady decrease in it from the equilibrium value, indicating softening of the solid with activity. As activity increases, the fraction of trajectories following linear response decreases to vanish near $\Lambda=10$; see Appendix-F. The decrease in $G$ with activity can be described by ${\cal G}={\cal G}_{0}-{\cal G}_{1}\Lambda^{2}$ with ${\cal G}_{0}=109.84$ and ${\cal G}_{1}=0.094$. To further explore this softening and the fluidization transition, we compute structural indicators, including solid order $\psi_{\mathbf{G}}$, hexatic order $\psi_{6}$, and defect fraction $\delta_{d}$.